For a solid cylinder of radius r and height h:

(i) write down the expression for the total surface area A,

(ii) write down the expression for the volume V,

(iii) using these expressions express V in terms of A and r in its simplest

form.


1
Expert's answer
2021-03-11T06:20:32-0500

(i)

"A=2\pi rh+2\pi r^2"

(ii)


"V=\pi r^2h"

(iii)

"h=\dfrac{A-2\pi r^2}{2\pi r}"

"V=\dfrac{r(A-2\pi r^2)}{2}"